(B):if 10g of phosphine is made to react with 15g of chlorine gas,calculate the percentage yield if 5.5g of phosphorus pentachloride was experimentally produced
(C)calculate the wavelength (in meters) associated with an 80g mass moving at 3km/hr
1
Expert's answer
2019-04-01T07:47:51-0400

B. PH3 + 4Cl2 → PCl5 + 3HCl

n(PH3) = 10g/34g/mol = 0.294 mol

n(Cl2) = 15g/71g/mol = 0.211 mol

PH3 is in excess, Cl2 is the limiting reactant.

n(PCl5) = 0.211/4 = 0.053 mol

m(PCl5) = 0.053 mol*208.5 g/mol = 11.05 g

Percentage yield of PCl5 = 5.5g/11.05g = 0.498 = 49.8%

C. Wavelength: Λ = h/mv = 6.63*10-34kg·m2/s / 0.08kg·0.833m/s = 9.95·10-33